# Marginal and conditional summary measures: transportability and compatibility

### v1

@@ 5 Discussion and concluding remarks
While ICH E9 (R1) lacks explicit guidelines for estimands in evidence synthesis, it already warns of potential estimand incompatibilities, cautioning against naïve comparisons across trials without consideration of the estimand that is addressed in each.
@@
We have clarified the interpretation of different marginal and conditional summary measures, and illustrated practical implications of estimand incompatibility for indirect treatment comparisons. Several novel covariate adjustment methodologies have been developed in this area, with different approaches targeting different summary measures. Nevertheless, the described estimand incompatibility issues may also apply to other areas of evidence synthesis, such as the naïve pooling of marginal and conditional estimates in pairwise – unadjusted or covariate-adjusted – meta-analyses of direct treatment comparisons (stat00513, 2016), or across indirect treatment comparisons or network meta-analyses that do not adjust for covariates (stat07909, 2014).
@@
Estimand incompatibility issues are particularly concerning for evidence synthesis based on published summary effect measures, as different trials may report different summary measures. Full access to subject-level data for all trials is rarely available in practice. Nevertheless, this would be beneficial to resolve estimand incompatibility issues by allowing for an IPD (network) meta-regression, the “gold standard” evidence synthesis approach that ML-NMR seeks to generalize. Such approach can: (1) produce any desired marginal or conditional summary measure without incompatibility issues; (2) resolve heterogeneity due to population differences; and (3) perform treatment comparisons in any given target population of interest.
@@
Many evidence syntheses of aggregate-level data are based on raw outcome data reported in each study arm, e.g., event counts and sample sizes, and thus combine marginal treatment effects across studies. However, purely prognostic covariates that are not conventionally described as effect modifiers – in the sense that they do not interact with treatment or induce treatment effect heterogeneity across subjects – can modify marginal treatment effects and compromise transportability, particularly for non-collapsible summary measures. This is generally under-appreciated in the evidence synthesis literature, with the exception of a recent article on network meta-analysis (Riley et al., 2023). Whether particular covariates can be classed as effect modifiers – thereby leading to departures from transportability, constancy or external validity – depends on the summary measure, and on whether it is marginal or conditional.
@@
While the summary measure is at the heart of the estimand framework, as described by ICH E9 (R1), it is not an explicit component of frameworks such as PICO (Population, Intervention, Comparator, Outcome), typically used for scoping in evidence synthesis, and often a secondary consideration when postulating research questions in evidence synthesis. Ultimately, we seek to underscore the importance of the summary measure as a fundamental component of any clinical research question.
@@
Within evidence synthesis, we can think of estimands at two different hierarchical levels: (1) at the trial level; and (2) at the meta-analytical level (Remiro-Azócar, 2022a). Both levels are important. There is a need to align estimands across different studies, as highlighted by this article. Nevertheless, there is also a need to align the estimand targeted at the meta-analytical level with a relevant clinical research question, motivating the shift towards more targeted meta-analyses that enhance applicability with respect to healthcare decision-making contexts (Ades et al., 2024; Remiro-Azócar & Gorst-Rasmussen, 2024).

### v3

@@ 5 Discussion and concluding remarks
While ICH E9 (R1) lacks explicit guidelines for estimands in evidence synthesis, it warns of potential estimand incompatibilities, cautioning against naïve comparisons across trials without considering the estimand that is addressed in each.
@@
We have clarified the interpretation of different marginal and conditional summary measures, and illustrated practical implications of estimand incompatibility for indirect treatment comparisons. Several novel covariate adjustment methodologies have been developed in this area, with different approaches targeting different summary measures. The estimand incompatibility issues described in this article may also apply to other areas of evidence synthesis, such as the naïve pooling of marginal and conditional estimates in pairwise – unadjusted or covariate-adjusted – meta-analyses of direct treatment comparisons (stat00513, 2016), and across indirect treatment comparisons or network meta-analyses that do not adjust for covariates (stat07909, 2014).
@@
Estimand incompatibility is particularly concerning for evidence synthesis based on published summary effect measures, as different trials may report different summary measures. Full access to subject-level data for all trials is rarely available in practice. However, this would be beneficial to resolve estimand incompatibility issues by allowing for an IPD (network) meta-regression, the “gold standard” evidence synthesis approach that ML-NMR seeks to generalize. Such an approach can produce any desired marginal or conditional summary measure without incompatibility issues, resolve heterogeneity due to population differences, and perform treatment comparisons in any given target population of interest.
@@
Many evidence syntheses of aggregate-level data are based on raw outcome data reported in each study arm, e.g., event counts and sample sizes, and thus combine marginal treatment effects across studies. However, purely prognostic covariates that are not conventionally described as effect modifiers – which do not interact with treatment nor induce treatment effect heterogeneity across subjects – can modify marginal treatment effects and compromise transportability, particularly for non-collapsible summary measures. This is generally under-appreciated in the evidence synthesis literature, with the exception of a recent article on network meta-analysis (Riley et al., 2023). Whether particular covariates can be classed as effect modifiers – thereby leading to departures from transportability, constancy or external validity – depends on the summary measure, and on whether it is marginal or conditional.
@@
While the summary measure is at the heart of the estimand framework, as described by ICH E9 (R1), it is not an explicit component of frameworks such as PICO (Population, Intervention, Comparator, Outcome), typically used for scoping in evidence synthesis, and often a secondary consideration when postulating research questions in evidence synthesis. We emphasize the importance of the summary measure as a fundamental component of any clinical research question.
@@
Within evidence synthesis, we can think of estimands at two different hierarchical levels: at the trial level and at the meta-analytical level (Remiro-Azócar, 2022a). Both levels are important. There is a need to align estimands across different studies, as highlighted by this article. There is also a need to align the estimand targeted at the meta-analytical level with a relevant clinical research question, motivating the shift towards more targeted meta-analyses that enhance applicability with respect to healthcare decision-making contexts (Ades et al., 2024; Remiro-Azócar & Gorst-Rasmussen, 2024).